1. Common Stock = 100,000. Accumulated Retained Earnings = 50,000. Total Assets = 400,000. Net income = 20,000. What is Return on Assets? (Express your answer as a percentage)

____________________

2. Assume Days Sales Outstanding is 30 days and Sales are 8,000,000. What are Accounts Receivable?

A. $800,000

B. $543,762

C. $657,534

D. $742,387

3. Assume the Profit Margin is 3%. Total Asset Turnover is 3x and the Equity Multiplier is 1.5. What is Return on Assets?

A. 13.5%

B. 3%

C. 9%

D. 1%

4. Assume the Profit Margin is 3%. Total Asset Turnover is 4x and the Equity Multiplier is 2. What is Return on Equity? (Express your answer as a percentage)

__________________

Thank you in advance!!!
Answer #1

1) Return on asset= Net income/total asset=20,000/400,000=5%

2) DSO= Accounts receivable/Sales*number of days in month

DSO=30

No of days =30

Sales=$800,000

30=AR/800,0000*30

Accounts receivables=$800,000

3)

ROA=Asset turnover*Profit margin

Return on asset= 3*3=9%

4)

ROE=Asset turnover*Profit margin*Equity multiplier

=4*3*2=24%